1/v+5=8 solve for v
Answer by Boreal(15235)

1/v+5=8
multiply everything by v
1+5v=8v
3v=1
v=(1/3)
check
1/(1/3)=3
3+5=8
============
If it is 1/(v+5)=8
8v+40=1
8v=39
v=-39/8